Circular No. 6180

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1995 M1 (SHOEMAKER-LEVY 4) J. V. Scotti, Lunar and Planetary Laboratory, reports his recovery

of comet P/1991 C2 (= 1990 XII = 1991f) with the 0.9-m Spacewatch telescope at Kitt Peak. The comet is of stellar appearance. The indicated correction (ignoring perturbations) to the prediction by S. Nakano on MPC 23482 is ΔT ∼ +0.6 day.

1995 UT  α_{2000}  δ_{2000}  m_{2}
June 22.45278  22^{h}31^{m}35.^{s}85  −7^{o}37′23′′.7  21.9
22.45800  22 31 35.94  −7 37 23.4  21.9
23.41216  22 31 40.18  −7 37 41.6  21.7
23.45819  22 31 40.37  −7 37 42.8

Orbital elements from 31 observations 1991–1995, mean residual 0′′.85: Epoch = 1996 Dec. 23.0 TT

T = 1997 Jan. 12.1950 TT  ω = 301.^{o}9866
e = 0.420477  Ω = 152.0973 }2000.0
q = 2.021244 AU  i = 8.4734

a = 3.487771 AU n^{o} = 0.1513150 P = 6.514 years

COMET P/1995 M2 (PARKER-HARTLEY) Scotti also reports his recovery of comet P/1989 E1 = 1986 TF (= 1989

XXXVI = 1989i), the second-night images being obtained by R. Jedicke. There was a 7′′ coma (m_{1} = 19.2) with a tail extending 0′.88 in p.a. 247^{o} on June 23 and 0′.76 in p.a. 248^{o} on June 25. The indicated correction to the prediction by Nakano on MPC 22031 is ΔT = +0.7 day.

1995 UT α_{2000} δ_{2000} m_{2} Observer
June 23.43116 0^{h}38^{m}25.^{s}69 +9^{o}17′58′′.1 20.5 Scotti
23.45167 0 38 26.58 +9 18 05.0  _{”}
25.43013 0 39 54.87 +9 29 04.5 20.2 Jedicke
25.43345 0 39 55.01 +9 29 05.6  _{”}

Orbital elements from 27 observations 1986–1995, mean residual 0′′.98: Epoch = 1996 June 6.0 TT

T = 1996 June 25.9910 TT  ω = 181.^{o}1934
e = 0.290368  Ω = 244.3013 }2000.0
q = 3.046185 AU  i = 5.1835
a = 4.292628 AU n^{o} = 0.1108203 P = 8.894 years

1995 June 28 Brian G. Marsden